Technical Observing - LBC IQ Tests 20140322 UT

Observers: JHill, OKuhn, MEdwards (Tucson - remote room)
Telescope Operator: SAllanson (LBTO)

Summary

First Half

We got decent binocular data for the LBC IQ tests. It has not yet been analyzed. This data was taken on the 2010J2 field with possible science value.

LUCI2/AGW2 Transform struggled with a long string of problems, and did not collect useful transform data.

We verified that the SX M2 swing arm and hexapod were in the correct position to close out Issue 5030. (All the pointing and collimation models need to be redone with the mass of LUCI2, MODS2 and ARGOS on the telescope.)

We verified that the star image is within 2+-1.5 arcsec of the nominal position when we switch to the ARGOS focal station and the dichroic. The 5 arcsec errors seen on Wednesday were the result of AGw probe centering and transformation problems at different rotator angles.

Second Half

ARGOS Laser Propagation Tests

Plan

2 hours of LBC IQ Tests

  • Preset to Pointing Star PA=0
    • DOFPIA to Collimate
    • Co-Point the two LBCs
    • Run OBs for in-focus LBC images
    • Run Superfoc for slightly defocussed LBC images
    • Run XXXX for full frame LBCFPIA pupil images
  • Apply coma-free pointing +40 arcsec Tip in the Y/RX direction and repeat
  • Apply coma-free pointing -40 arcsec Tip in the Y/RX direction and repeat
  • Apply coma-free pointing -40 arcsec Tilt in the X/RY direction and repeat
  • Choose the best of the above positions and change PA to 180 deg

Details

Open

01:30 Open to a mostly clear sky with scattered thin cirrus

LBC IQ tests to evaluate binodal astigmatism

(John driving PSF and DOFPIA, Olga driving LBC)

Christian's image analysis for the ellipticity is at: http://abell.as.arizona.edu/~lbtsci/cv/lbciq20140322.html

01:54 TRACK preset for LBCs

01:57 DOFPIA although sky is still bright

02:09 TRACK preset to 2010J2

02:15 DOFPIA with actual pupils

02:21 Play 2010J2 OB just past meridian at EL=80.4 deg. Guiding on Blue looks peculiar in the EL direction. Blue Science Chip is saturated in 3 min.

02:30 Play modified 2010J2 OB with two 60s Blue and one 120 Red exposure. Cirrus might be 1 mag thick.

02:47 DOFPIA

02:52 Play superfoc_coarse (PA=0)

02:58 Play RB_RV_exin4

03:01 Play 2010J2 OB (for 2-3 dithers) why did Blue stop guiding after the first 2?

Was Blue already at Tip 0 Tilt -40? Yes - maybe that explains the peculiar guiding.

03:12 Change Blue to Tip=0 Tilt=0 and Red to Tip=0 Tilt=-40 (Coma-free Pointing sent to M1 by PSF)

03:13 DOFPIA

03:17 Play superfoc_coarse

03:21 Play RB_RV_exin4

03:25 Play 2010J2 OB (for 2-3 dithers) -- Stopped this OB because Blue tech chips weren't guiding.

Power cycle the Blue Tech Chips. (as root with the power gui)

03:35 But things were still confused, so do a full power cycle and restart of the lbc software

03:47 Play 2010J2 OB but not connected to telescope. That's why the images at 034413/18 are rotationally trailed about the rotator center.

DIMM is varying between 1.0 and 1.4 arcsec.

03:51 Play 2010J2 OB 120 s exposures for Blue and Red - sky is approaching photometric.

04:00 Move to Tip=+40 Tilt=0 coma-free pointing on both sides

04:01 DOFPIA

04:05 Play superfoc_coarse

04:09 Play RB_RV_exin4

04:12 Play 2010J2 OB (3 x 120 sec images on each side)

04:26 Reconfigure to LUCI at Front Bent on both sides.

Authorize LUCI@RFBG

AGW2 transform measurements

(Michelle driving LUCI2 and IDL, Steve driving IRAF, John assisting with AGw )

Pointing mask is inserted. Filter is Brackett Gamma.

Script data in /home/lbto/data/20140322/rfbg/

04:45 Preset to a Pointing Star IE=-10 CA=-191 at EL=47 deg. We are looking directly into the wind at 12 m/sec.

AdSec DX RIP........

Michelle having problems with LUCI2 control panel.

04:51 ACTIVE Preset -- seeing on RFBG guider is 0.9-1.2 arcsec.

05:02 ACTIVE Preset to Stone_E1_15 on-axis

img 0003 is dark frame (or back of probe)

05:07 Stop Guiding and retract probe img 0004

LUCI2 is stuck in countdown

From: Michelle Edwards 
Date: Fri, 21 Mar 2014 23:11:19 -0700

Took an image through GEIRS and had no initial problems, but while trying to
take images through the readout LUCI2 pink engineering screen, saw no countdown
of the exposure time. Called D. Thompson who led me through a series of
restarts in the engineering panel. (readout manager, etc) which initially
seemed to fix the problem, however the software hung again and I did the
restart again. Things seemed ok, and we started to try to take the initial
on-axis data when we then had a crash of GEIRS (evident in a very negative
countdown of the readout manager). Restarted GEIRS.

Realized then that something was not right with the data dictionary; call to T.
Sargent uncovers that the telescope service and data dictionary almost
certainly should have been restarted. Restart is successful from engineering
panel.

Now LUCI is happy, telescope service is happy. Try to start transform again and
unfortunately the transform program fails -- looking for image
ght_wfscimage000048.fits which should have been right_wfscimage000048.fits.
Kills program, cannot proceed, which is fine since the star is already setting.

05:18 ACTIVE Preset to Stone_E1_15 on-axis

05:23 Stop Guiding and retract probe

LFBG Coma Check

05:28 Authorize LUCI on both sides

05:33 ACQUIRE preset on-axis parallactic on both sides (just a bit of coma and astigmatism, but easily capturable by Active Optics)

DX AdSec RIPs because anemometer process died.

05:43 PseudoMono on Right, ACTIVE preset PA=0

GCSL waits for right rotator to unwrap

05:47 Stop guiding, retract probe

05:49 ACTIVE preset PA=-135

05:51 Stop guiding, retract probe

05:53 transform_collect (failed because of a string problem with the name "ght_wfsc.......fits")

RFBG Dichroic Pointing Tests

(John directing, Steve driving, Marco moving dichroic)

RFBG focal station with dichroic removed

Authorize LUCI @ Front Bent on both sides

still in PMMode on right side

06:03 ACQUIRE parallactic on-axis BS9143, Center the star IE=-6 CA=-190

Shell RIPs w/TSS

06:08 ACQUIRE parallactic on-axis BS9143, Center the star IE=-6 CA=-191

06:11 ACQUIRE on-axis PA=90 x=2.23 y=612.5 (rotangle=84.8), Center the star IE=-10 CA=-195

06:15 ACQUIRE parallactic on-axis x=2.41 y=612.37 BS9143 rotangle=3.4, Center the star IE=-7 CA=-192

ARGOS focal station with dichroic inserted

06:19 Authorize LUCI @ ARGOS on both sides (hit DX M2 limit on the way ????)

06:22 ACQUIRE parallactic on-axis BS9143, Center the star IE=-9 CA=-193

Passive preset on left side didn't go through. Apparently ARGOS is not authorized for PMMode.

06:26 Spin rotator by 360 - makes a 5 arcsec diameter circle to the left side.

RFBG focal station with dichroic removed

06:30 Authorize LUCI at front bent, and remove dichroic

06:33 ACQUIRE parallactic on-axis BS9143

Center the star IE=-8 CA=-193 (now different by only 1 arcsec) sigma of the measurement is 1.5 arcsec.

ARGOS Laser Operations

ARGOS focal station with dichroic inserted

06:36 Authorize LUCI@ARGOS on both sides and put dichroic in

06:38 ACQUIRE preset, center the star at IE=-7 CA=-193

06:40 ACTIVE preset collimating

06:45 Slew to AZ=120 EL=90 - some wispy cirrus now in the north, and moon rising in the south east

(these operations are not logged here)

-- %USERSIG{JohnHill - 2014-03-22}%

Comments

 
Topic revision: r3 - 28 Mar 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback